# Absolutely Maximal Contextual Correlations

### v1

@@ Abstract
The foundational work by Bell led to an interest in understanding non-local correlations that arise from entangled states shared between distinct, spacelike-separated parties, which formed a foundation for the theory of quantum information processing. We investigate the question of maximal correlations analogous to the maximally entangled states defined in the entanglement theory of multipartite systems. To formalize this, we employ the sheaf-theoretic framework for contextuality, which generalizes non-locality. This provides a new metric for correlations called contextual fraction (CF), which ranges from 0 (non-contextual) to 1 (maximally contextual). Using this, we have defined the absolutely maximal contextual correlations (AMCC), which are maximally contextual and have maximal marginals, which captures the notion of absolutely maximal entangled (AME) states. The Popescu-Rohrlich (PR) box serves as the bipartite example, and we construct various extensions of such correlations in the tripartite case. An infinite family of various forms of AMCC is constructed using the parity check method and the constraint satisfiability problem (CSP) scheme. We also demonstrate the existence of maximally contextual correlations, which do not exhibit maximal marginals, and refer to them as non-AMCC. The results are further applied to secret sharing and randomness extraction using AMCC correlations.
@@ I Introduction
The correlations obtained by measuring an entangled state constitute another important form of nonclassicality, both from a resource-theoretic and an applied perspective. In a Bell experiment, these correlations can be viewed abstractly as a “black box” characterized by inputs and outputs, with its internal physical structure being irrelevant. This abstraction makes them a powerful resource for device-independent protocols Pironio et al. (2010); Moreno et al. (2020). Correlations that violate a Bell inequality are termed nonlocal and belong to a strictly larger set than those admitting a local hidden-variable model. Tsirelson established that the quantum maximum for the paradigmatic Clauser-Horne-Shimony-Holt (CHSH) expression is $2\sqrt{2}$ Cirel’son (1980). More generally, post-quantum nonlocal correlations—supersets of the quantum set—have also been studied, as exemplified by the Popescu-Rohrlich box Popescu and Rohrlich (1994).
@@
The theories of entanglement and nonlocal correlations are deeply intertwined, presenting significant conceptual intricacies Schmid et al. (2023). This interplay is clarified by the framework of local operations and shared randomness (LOSR) Schmid et al. (2020); Rosset et al. (2020), which serves as the correlational analog to LOCC in entanglement theory. To understand genuine multipartite correlations, a more general operational approach employs wiring and classical communication prior to inputs (WCCPI) Gallego et al. (2012), demonstrating that earlier definitions were incomplete Bancal et al. (2011). A fundamental difficulty arises from a key divergence: while the classification of pure entangled states into equivalence classes under LOCC and SLOCC is well-established, no analogous structure exists for nonlocal correlations. In the bipartite case, LOCC yields a single equivalence class of maximally entangled states—a definitive “golden unit.” However, no corresponding unique unit has been identified for correlations. Concepts such as single-copy or asymptotic distillation do not translate directly from entanglement theory, preventing a straightforward mapping from maximal entanglement to maximal correlation. Consequently, many correlation sets are incomparable in the asymptotic limit Ghosh et al. (2024). Thus, a central challenge is to formulate rigorous definitions for maximal entanglement and maximal correlations in genuine multipartite systems, an area that remains actively unresolved.
@@
Correlations from an experimental setup can be viewed more generally using a sheaf-theoretic approach Abramsky and Brandenburger (2011). Any dataset described by $(n,m,o)$, where $n$ is the number of parties, $m$ is the number of measurements per party, and $o$ is the set of possible outcomes, can be formalized as a measurement scenario. In this framework, compatible sets of measurements (called contexts) and their possible outcomes are organized so that probability distributions over outcomes define the empirical model.
A no-signaling empirical model is one in which outcome probabilities do not depend on which other measurements are performed, generalized by the compatibility condition. If these probabilities can be derived from a single global assignment, the model is noncontextual; otherwise, it is contextual. More formally, any empirical model can be represented as a linear system: if this system admits a non-negative solution, the model is noncontextual; otherwise, it is contextual. In the possibilistic case, a stronger criterion defines the notion of a strongly contextual model. Any empirical model can be expressed as a convex combination of a noncontextual model (with coefficient non-contextual fraction, $\mathrm{NCF}=p$) and a strongly contextual model (with coefficient contextual fraction $\mathrm{CF}=1-p$). If CF=1 ($p=0$), the model is maximally contextual; if CF=0 ($p=1$), the model is noncontextual Abramsky et al. (2017). It has been shown that a model is strongly contextual if and only if it is maximally contextual Abramsky and Brandenburger (2011).
@@
These results and framework has broader implications for our understanding of computational advantage. In particular, within quantum computation, this advantage manifests as the ability to transition from limited computational power, characterized by the parity complexity class, to full quantum universality Anders and Browne (2009). This transition has been demonstrated using GHZ-type correlations and Mermin-type arguments. Moreover, the contextuality has been interpreted in various ways Raussendorf (2013); Okay et al. (2017); Abramsky et al. (2012); Raussendorf (2019); Frembs et al. (2018); Walleghem et al. (2024); Raussendorf et al. (2023); Okay and Raussendorf (2020); Xu et al. (2020); Saha et al. (2019); Beer and Osborne (2018).
@@
It is evident that contextuality is a more general phenomenon than nonlocality, and that correlations exhibiting maximal contextuality are termed maximal contextual correlations, as they do not contain any local fraction. To extend the idea of genuine multipartite entanglement, a special class of entanglement is defined as absolutely maximally entangled (AME) states Helwig et al. (2012). These pure states are maximally entangled, such that for any bipartite reduction, they yield maximally mixed states. In this back drop, we would like to ask the question what is the analogues of AME in the theory of correlations. In this work, we extend this to genuine multipartite correlations and define a new class of correlations as absolutely maximal contextual correlations (AMCC). These correlations represent the maximal contextual correlations, which are also maximally random marginals. In the $(2,2,2)$ scenario, there are exactly eight such AMCCs which are Popescu-Rohrich (PR) boxes. However, in the multipartite scenario like $(3,2,2)$, there are infinitely many such correlations that qualify as AMCCs and reside on the face of the polytope. The mathematical construction of these correlations can be accomplished using the parity check method or the constraint satisfiability problem (CSP) approach. Apart from AMCCs, there exists another class of correlations that are maximally contextual but not maximally random marginals; these are constructed exclusively using the CSP method. The AMCCs correspond to correlations that are analogous to absolutely maximally entangled (AME) states. This property is of major significance in device-independent randomness extraction due to the maximal marginal property, and such correlations can also be utilized in secure communication protocols.
@@ II Formalism
for all $j$, all outcomes $x_{1},\dots,x_{j-1},x_{j+1},\dots,x_{n}$, and all settings $X_{1},\dots,X_{j-1},X_{j},X_{j}^{\prime},X_{j+1},\dots,X_{n}$. The set of probability distributions satisfying these constraints forms a convex polytope known as the no-signaling polytope, with dimension
@@
$p(x_{1},\dots,x_{n}|X_{1},\dots,X_{n})=\int d\lambda\,\rho(\lambda)\,p(x_{1}|X_{1},\lambda)\cdots p(x_{n}|X_{n},\lambda),$ (2)
where $\rho(\lambda)\geq 0$, $\int d\lambda\,\rho(\lambda)=1$, and $p(x_{i}|X_{i},\lambda)$ denotes the probability of obtaining outcome $x_{i}$ for measurement $X_{i}$ given the hidden variable $\lambda$. Here $\lambda$ represents a set of pre-assigned values for all possible measurement outcomes. The corresponding deterministic probability vectors are defined as
$p^{\lambda}(x_{1},\dots,x_{n}|X_{1},\dots,X_{n})=\begin{cases}1,&\text{if }\forall i,\;\lambda_{iX_{i}}=x_{i},\;i\in\{1,\dots,n\},\\ 0,&\text{otherwise}.\end{cases}$ (3)
The set of local correlations $P$ is the convex hull of these deterministic points, forming the local polytope
@@
For the Bell polytope $P\subseteq\mathbb{R}^{r}$, consider an inequality $\mathbf{a}\cdot\mathbf{p}\geq a_{0}$, where $(\mathbf{a},a_{0})\in\mathbb{R}^{r+1}$. This inequality is called a valid Bell inequality for $P$ if it holds for all $\mathbf{p}\in P$. To verify validity, it suffices to check that it is satisfied by all extreme points of $P$, i.e., by the deterministic vectors. Given such an inequality, the corresponding face of the polytope is defined by the equality condition $F=\{\mathbf{p}\in P\;|\;\mathbf{a}\cdot\mathbf{p}=a_{0}\}$. A face with dimension exactly one less than that of the full polytope is called a facet Pironio (2005). Thus, for any proper face $F$ (with $F\neq\varnothing$ and $F\neq P$), the dimension satisfies $\dim F\leq\dim P-1$. These proper faces, which also satisfy the no-signaling and normalization conditions, are defined by zero constraints:
@@
where $p$ is an element of $\mathbf{p}=\{p(x_{1},\cdots,x_{i}\mid X_{1},\cdots,X_{i})_{i\leq\dim P-1}\}$, and $\mathbf{NS}\subset\mathbb{R}^{r}$ denotes the set of no-signaling correlations. By imposing these zero constraints, the correlations are brought to their maximally reduced form, corresponding to the lowest-dimensional face. The resulting face correlations yield a set of deterministic constraints that cannot be satisfied simultaneously by any global assignment. In other words, this construction leads directly to AVN (all-versus-nothing) proofs (see the Appendix(A)). Hence, all the zeros define a face of the no-signaling polytope, and every deterministic local vertex $\mathbf{p}^{L}\in P$ violates at least one zero constraint. This implies that $F^{NS}\cap\mathbf{p}^{L}=\phi$; in other words, the correlations on the face $F^{NS}$ in Eq.(4) do not contain any local components Liu et al. (2024). These face correlations are referred to as maximal correlations, which are discussed in the next section.
@@ II.1 Contextual correlations
The correlations corresponding to maximal contextuality lie on the face of the nonlocal polytope, as they do not contain any local fractions  [^1] Any empirical model can be decomposed as a convex mixture of contextual and noncontextual models, with contextuality being a more general concept than nonlocality. The convex decomposition of a nonlocal correlation can be expressed as $p(x_{1},\ldots,x_{n})=q^{L}p_{L}(x_{1},\ldots,x_{n})+q^{NL}p_{NL}(x_{1},\ldots,x_{n}),$ where (pL) denotes the local part and (pNL) the nonlocal part. This decomposition is a special case within the broader framework of contextuality. Hence, the condition that the contextual fraction $CF=1$ corresponds to a special case where the nonlocal fraction is also maximal. This is significant because such correlations also correspond to all-versus-nothing (AVN) proofs, such as Mermin’s arguments Liu et al. (2024). In the probabilistic form of AVN, the presence of zeros determines the structure of these correlations; these zeros specify the facets that maximally reduce the dimension of the face while still satisfying the no-signaling and normalization conditions. We utilize this property in Section III for construction purposes. While the concept of maximal entanglement is well established for quantum states, its direct analogue in terms of correlations remains somewhat ambiguous. To clarify the notion of maximality in correlations, we take absolutely maximally entangled (AME) states Helwig et al. (2012) as a reference point, since their genuine bipartite marginals are maximally mixed. Accordingly, we focus on correlations that are not only maximally contextual but also exhibit maximal marginality. This approach leads to the identification of a new class of correlations and provides a precise definition of maximality in the context of correlations, which will be introduced in the next section.
@@ III Absolutely maximal contextual correlations
Furthermore, the mathematical formalism allows for the simulation of empirical models and their associated measurement scenarios, as defined in Section II. In this section, we utilize these concepts to introduce a new class of maximal contextual correlations, based on the maximally contextual model. These correlations are of particular interest as they satisfy the all-versus-nothing (AVN) proof, and are therefore referred to as face-non-signaling (FNS) correlations Liu et al. (2024). The required mathematical definitions are defined below in this Section.
@@
$e|_{U}=p(x_{1},\dots,x_{k}\mid X_{1},\dots,X_{k})=\frac{1}{2^{k}}\quad\forall\;k<n,$ (16)
@@
$P_{\text{PR}}^{\alpha\beta\gamma}(x_{1}x_{2}|X_{1}X_{2})=\begin{cases}\frac{1}{2},&\begin{aligned} x_{1}\oplus x_{2}&=X_{1}X_{2}\oplus\alpha X_{1}\\ &\oplus\beta X_{2}\oplus\gamma,\end{aligned}\\ 0,&\text{otherwise},\end{cases}$ (17)
@@
$P_{\text{PR}}^{\alpha\beta\gamma}(x_{1}x_{2}|X_{1}X_{2})=\frac{1}{2}\left(\delta_{0}^{x_{1}}\delta_{f}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{f\oplus 1}^{x_{2}}\right),$ (18)
@@
$\begin{split}&P(x_{1}|X_{1})=\frac{1}{2}\left(\delta_{0}^{x_{1}}+\delta_{1}^{x_{1}}\right),\\ &P(x_{2}|X_{2})=\frac{1}{2}\left(\delta_{0}^{x_{2}}+\delta_{1}^{x_{2}}\right)\end{split}$ (19)

### v2

@@ Abstract
The foundational work by Bell led to an interest in understanding non-local correlations that arise from entangled states shared between distinct, spacelike-separated parties, which formed a foundation for the theory of quantum information processing. We investigate the question of maximal correlations analogous to the maximally entangled states defined in the entanglement theory of multipartite systems. In this work, we define the maximality of nonlocal correlation as being analogous to the absolutely maximally entangled state. To formalize this, we employ the sheaf-theoretic framework for contextuality, which generalizes non-locality. This provides a new metric for correlations called contextual fraction (CF), which ranges from 0 (non-contextual) to $1$ (maximally contextual). Using this, we have defined the absolutely maximal contextual correlations (AMCC), which are maximally contextual and have maximal marginals. The Popescu-Rohrlich (PR) box serves as the bipartite example, and we construct various extensions of such correlations in the tripartite case. An infinite family of various forms of AMCC is constructed using the parity check method and the constraint satisfiability problem (CSP) scheme. We also demonstrate the existence of maximally contextual correlations, which do not exhibit maximal marginals, and refer to them as non-AMCC. The results are further applied to secret sharing and randomness extraction using AMCC correlations.
@@ I Introduction
The correlations obtained by measuring an entangled state constitute another important form of nonclassicality, both from a resource-theoretic and an applied perspective. In a Bell experiment, these correlations can be viewed abstractly as a “black box” characterized by inputs and outputs, with its internal physical structure being irrelevant. This abstraction makes them a powerful resource for device-independent protocols Pironio et al. (2010); Moreno et al. (2020). Correlations that violate a Bell inequality are termed nonlocal and belong to a strictly larger set than those admitting a local hidden-variable model Rai et al. (2019). Tsirelson established that the quantum maximum for the paradigmatic Clauser-Horne-Shimony-Holt (CHSH) expression is $2\sqrt{2}$ Cirel’son (1980). More generally, post-quantum nonlocal correlations—supersets of the quantum set—have also been studied, as exemplified by the Popescu-Rohrlich box Popescu and Rohrlich (1994).
@@
The theories of entanglement and nonlocal correlations are deeply intertwined, presenting significant conceptual intricacies Schmid et al. (2023). This interplay is clarified by the framework of local operations and shared randomness (LOSR) Schmid et al. (2020); Rosset et al. (2020), which serves as the correlational analog to LOCC in entanglement theory. To understand genuine multipartite correlations, a more general operational approach employs wiring and classical communication prior to inputs (WCCPI) Gallego et al. (2012); Dutta et al. (2020), demonstrating that earlier definitions were incomplete Bancal et al. (2011). A fundamental difficulty arises from a key divergence: while the classification of pure entangled states into equivalence classes under LOCC and SLOCC is well-established, no analogous structure exists for nonlocal correlations. In the bipartite case, LOCC yields a single equivalence class of maximally entangled states—a definitive “golden unit.” However, no corresponding unique unit has been identified for correlations. Concepts such as single-copy or asymptotic distillation do not translate directly from entanglement theory, preventing a straightforward mapping from maximal entanglement to maximal correlation. Consequently, many correlation sets are incomparable in the asymptotic limit Ghosh et al. (2024). Thus, a central challenge is to formulate rigorous definitions for maximal entanglement and maximal correlations in genuine multipartite systems, an area that remains actively unresolved.
@@
Correlations from an experimental setup can be viewed more generally using a sheaf-theoretic approach Abramsky and Brandenburger (2011). Any dataset described by $(n,m,o)$, where $n$ is the number of parties, $m$ is the number of measurements per party, and $o$ is the set of possible outcomes, can be formalized as a measurement scenario. In this framework, compatible sets of measurements (called contexts) and their possible outcomes are organized so that probability distributions over outcomes define the empirical model. A no-signaling empirical model is one in which outcome probabilities do not depend on which other measurements are performed, generalized by the compatibility condition. If these probabilities can be derived from a single global assignment, the model is noncontextual; otherwise, it is contextual.
Contextuality is a nonclassical feature indicating the impossibility of assignment to measurements independent of the measurement context. The KS theorem Kochen and Specker (2011), in this regard, provides a foundational understanding of such logical proofs. The Mermin square proofs Mermin (1990) with $9$ observables and the 18-ray vector Cabello et al. (1996), which contains the first minimal set of projectors after the 117-vector Kochen and Specker (2011), followed by the 13-ray vector Yu and Oh (2012) proof, show examples of state-independent KS theorem proofs. Fine’s theorem Fine (1982) is one of the pieces of evidence that leads us to the equivalence between the violation of the Bell inequality (CHSH) Clauser et al. (1969) and the impossibility of a global joint probability distribution. The KCBS inequality Klyachko et al. (2008) provides one of the simplest state-dependent proofs for a qutrit system, demonstrating that spatial separation is not required and that contextuality can be exhibited in a single system. In this respect, Bell nonlocality can be viewed as a special case of contextuality, which requires extra constraints such as specially separated parties and no-signaling conditions. The sheaf-theoretic approach makes this point more prominent and provides a general mathematical tool. Within this framework, any empirical model can be represented as a linear system: if this system admits a non-negative solution, the model is noncontextual; otherwise, it is contextual. In the possibilistic case, a stronger criterion defines the notion of a strongly contextual model. Any empirical model can be expressed as a convex combination of a noncontextual model (with coefficient noncontextual fraction, $\mathrm{NCF}=p$) and a strongly contextual model (with coefficient contextual fraction $\mathrm{CF}=1-p$). If $\mathrm{CF}=1$ ($p=0$), the model is maximally contextual; if $\mathrm{CF}=0$ ($p=1$), the model is noncontextual Abramsky et al. (2017). It has been shown that a model is strongly contextual if and only if it is maximally contextual Abramsky and Brandenburger (2011).
@@
These results and the framework have broader implications for our understanding of computational advantage. In particular, in quantum computation, contextuality enables transitions from limited computational power, characterized by the parity complexity class, to full quantum universality Anders and Browne (2009). This transition has been demonstrated using GHZ-type correlations and Mermin-type arguments, and the computational advantage of the measurement-based quantum computation (MBQC) model has been understood Raussendorf (2013); Abramsky et al. (2017). Contextuality also admits various mathematical characterizations, such as cohomological or topological obstruction Okay et al. (2017); Abramsky et al. (2012); Raussendorf (2019); Frembs et al. (2018); Walleghem et al. (2024); Raussendorf et al. (2023); Okay and Raussendorf (2020); Xu et al. (2020); Saha et al. (2019). Such a global obstruction can be geometrically visible in the bundle diagram Beer and Osborne (2018).
Although contextuality is an abstract concept, its experimental validation and implications have been extensively studied, including Mermin’s square and the KCBS inequality Kirchmair et al. (2009); Li et al. (2017). Quantifying and extending contextuality in quantum many-body systems is being investigated, which has motivated the translation of contextuality into a nonlocal game. Contextuality undermines the classical strategy Zhao et al. (2025); Hart et al. (2025). This also implies the connection between the phase of the quantum many-body system and contextuality. Recently, the Google research team demonstrated the implications of quantum contextuality for classical strategies with limited resources on a real-time quantum processor. That also shows the scalability of contextuality and operational relevance in the multiparty system et al. (2025). Contextuality has also been shown to certify intrinsic randomness, and its scalability is understood from a graph-theoretic perspective, which is experimentally feasible Liu and Ramanathan (2025). All the above instances involve parity conditions, which exhibit the importance of maximal contextuality.
@@
It is evident that contextuality is a more general phenomenon than nonlocality, and that correlations exhibiting maximal contextuality are termed maximal contextual correlations, as they do not contain any local fraction Liu et al. (2024); Rai et al. (2019). To extend the idea of genuine multipartite entanglement, a special class of entanglement is defined as absolutely maximally entangled (AME) states Helwig et al. (2012). These pure states are maximally entangled, such that for any bipartite reduction, they yield maximally mixed states. In this back drop, we would like to ask the question what is the analogues of AME in the theory of correlations. In this work, we extend this to genuine multipartite correlations and define a new class of correlations as absolutely maximal contextual correlations (AMCC). These correlations represent the maximal contextual correlations, which are also maximally random marginals. In the $(2,2,2)$ scenario, there are exactly eight such AMCCs, which are Popescu-Rohrich (PR) boxes. However, in the multipartite scenario like $(3,2,2)$, there are infinitely many such correlations that qualify as AMCCs and reside on the face of the polytope. The mathematical construction of these correlations can be accomplished using the parity check method or the constraint satisfiability problem (CSP) approach. Apart from AMCCs, there exists another class of correlations that are maximally contextual but not maximally random marginals; these are constructed exclusively using the CSP method. The AMCCs correspond to correlations that are analogous to absolutely maximally entangled (AME) states. This property is of major significance in device-independent randomness extraction due to the maximal marginal property, and such correlations can also be utilized in secure communication protocols.
@@ II Formalism
for all $j$, all outcomes $x_{1},\dots,x_{j-1},x_{j+1},\dots,x_{n}$, and all settings $X_{1},\dots,X_{j-1},X_{j},X_{j}^{\prime},X_{j+1},\dots,X_{n}$. The set of probability distributions satisfying these constraints forms a convex polytope known as the no-signaling polytope Pironio (2005); Rai et al. (2019), with dimension
@@
$p(x_{1},\dots,x_{n}\mid X_{1},\dots,X_{n})=\int d\lambda\,\rho(\lambda)\,p(x_{1}\mid X_{1},\lambda)\cdots p(x_{n}\mid X_{n},\lambda),$ (2)
where $\rho(\lambda)\geq 0$, $\int d\lambda\,\rho(\lambda)=1$, and $p(x_{i}\mid X_{i},\lambda)$ denotes the probability of obtaining outcome $x_{i}$ for measurement $X_{i}$ given the hidden variable $\lambda$. Here $\lambda$ represents a set of pre-assigned values for all possible measurement outcomes. The corresponding deterministic probability vectors are defined as
$p^{\lambda}(x_{1},\dots,x_{n}\mid X_{1},\dots,X_{n})=\begin{cases}1,&\text{if}\quad\forall i,\;\lambda_{iX_{i}}=x_{i},\;i\in\{1,\dots,n\},\\ 0,&\text{otherwise}.\end{cases}$ (3)
The set of local correlations $P$ is the convex hull of these deterministic points, forming the local polytope Pironio (2005); Rai et al. (2019)
@@
For the Bell polytope $P\subseteq\mathbb{R}^{r}$, consider an inequality $\mathbf{a}\cdot\mathbf{p}\geq a_{0}$, where $(\mathbf{a},a_{0})\in\mathbb{R}^{r+1}$. This inequality is called a valid Bell inequality for $P$ if it holds for all $\mathbf{p}\in P$. To verify validity, it suffices to check that it is satisfied by all extreme points of $P$, i.e., by the deterministic vectors. Given such an inequality, the corresponding face of the polytope is defined by the equality condition $F=\{\mathbf{p}\in P\;|\;\mathbf{a}\cdot\mathbf{p}=a_{0}\}$. A face with dimension exactly one less than that of the full polytope is called a facet Pironio (2005); Rai et al. (2019). Thus, for any proper face $F$ (with $F\neq\varnothing$ and $F\neq P$), the dimension satisfies $\dim F\leq\dim P-1$. These proper faces, which also satisfy the no-signaling and normalization conditions, are defined by zero constraints:
@@
where $p$ is an element of $\mathbf{p}=\{p(x_{1},\cdots,x_{i}\mid X_{1},\cdots,X_{i})_{i\leq\dim P-1}\}$, and $\mathbf{NS}\subset\mathbb{R}^{r}$ denotes the set of no-signaling correlations. By imposing these zero constraints, the correlations are brought to their maximally reduced form, corresponding to the lowest-dimensional face. The resulting face correlations yield a set of deterministic constraints that cannot be satisfied simultaneously by any global assignment. In other words, this construction leads directly to AVN (all-versus-nothing) proofs (see the Appendix (A)). Hence, all the zeros define a face of the no-signaling polytope, and every deterministic local vertex $\mathbf{p}^{L}\in P$ violates at least one zero constraint. This implies that $F^{NS}\cap\mathbf{p}^{L}=\phi$; in other words, the correlations on the face $F^{NS}$ in Eq. (4) do not contain any local components Liu et al. (2024); Rai et al. (2019). These face correlations are referred to as maximal correlations, which are discussed in the next section.
@@ II.1 Contextual correlations
The correlations corresponding to maximal contextuality lie on the face of the nonlocal polytope, as they do not contain any local fractions  [^1] Any empirical model can be decomposed as a convex mixture of contextual and noncontextual models, with contextuality being a more general concept than nonlocality. The convex decomposition of a nonlocal correlation can be expressed as $p(x_{1},\ldots,x_{n})=q^{L}p_{L}(x_{1},\ldots,x_{n})+q^{NL}p_{NL}(x_{1},\ldots,x_{n}),$ where (pL) denotes the local part and (pNL) the nonlocal part. This decomposition is a special case within the broader framework of contextuality. Hence, the condition that the contextual fraction $CF=1$ corresponds to a special case where the nonlocal fraction is also maximal. This is significant because such correlations also correspond to all-versus-nothing (AVN) proofs, such as Mermin’s arguments Liu et al. (2024). In the probabilistic form of AVN, the presence of zeros determines the structure of these correlations; these zeros specify the facets that maximally reduce the dimension of the face while still satisfying the no-signaling and normalization conditions. We utilize this property in Section (III) for construction purposes. While the concept of maximal entanglement is well established for quantum states, its direct analogue in terms of correlations remains somewhat ambiguous. To clarify the notion of maximality in correlations, we take absolutely maximally entangled (AME) states Helwig et al. (2012) as a reference point, since their genuine bipartite marginals are maximally mixed. Accordingly, we focus on correlations that are not only maximally contextual but also exhibit maximal marginality. This approach leads to the identification of a new class of correlations and provides a precise definition of maximality in the context of correlations, which will be introduced in the next section.
@@ III Absolutely maximal contextual correlations
Furthermore, the mathematical formalism allows for the simulation of empirical models and their associated measurement scenarios, as defined in Section (II). In this section, we utilize these concepts to introduce a new class of maximal contextual correlations, based on the maximally contextual model. These correlations are of particular interest as they satisfy the all-versus-nothing (AVN) proof, and are therefore referred to as face-non-signaling (FNS) correlations Liu et al. (2024). The required mathematical definitions are defined below in this Section.
@@
$e\mid_{U}=p(x_{1},\dots,x_{k}\mid X_{1},\dots,X_{k})=\frac{1}{2^{k}}\quad\forall\;k<n,$ (16)
@@
$P_{\text{PR}}^{\alpha\beta\gamma}(x_{1}x_{2}\mid X_{1}X_{2})=\begin{cases}\frac{1}{2},&\begin{aligned} x_{1}\oplus x_{2}&=X_{1}X_{2}\oplus\alpha X_{1}\\ &\oplus\beta X_{2}\oplus\gamma,\end{aligned}\\ 0,&\text{otherwise},\end{cases}$ (17)
@@
$P_{\text{PR}}^{\alpha\beta\gamma}(x_{1}x_{2}\mid X_{1}X_{2})=\frac{1}{2}\left(\delta_{0}^{x_{1}}\delta_{f}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{f\oplus 1}^{x_{2}}\right),$ (18)
@@
$\begin{split}&P(x_{1}\mid X_{1})=\frac{1}{2}\left(\delta_{0}^{x_{1}}+\delta_{1}^{x_{1}}\right),\\ &P(x_{2}\mid X_{2})=\frac{1}{2}\left(\delta_{0}^{x_{2}}+\delta_{1}^{x_{2}}\right)\end{split}$ (19)